OpenStudy (hippoharry):

Which of the binomials below is a factor of this trinomial? x2 + 4x - 60 A. x - 10 B. x2 - 6 C. x + 6 D. x - 6

OpenStudy (sweetburger):

This factors out to (x+10)(x-6)

OpenStudy (hippoharry):

im confused. what do i do after with (x+10)(x-6)

OpenStudy (hippoharry):

@sweetburger

OpenStudy (anonymous):

1*-60=-60 10-6=4 10*-6=-60

OpenStudy (anonymous):

\[x^2+(10-6)x-60=x^2+10x-6x-60=x(...+...)-6(...+...)\]